# Quantum Optimization Benchmark Library The Intractable Decathlon

### v1

@@ 4.6.3 Instances
Table 5 (left) outlines the parameters that were utilized in our experiments. The choice was based on established literature [87, 52, 217]. Using these parameters, we generated 4 instances each (1 containing the original data and 3 with randomly perturbed data) for 10, 50, 200, and 400 possible assets with two different time horizons: 10 days and 15 days. We used real S&P 500 data from January 1, 2024, to May 31, 202,4 for the asset prices $p_{it}$ and the corresponding covariances $\sigma_{ijt}$. The size of the instances is defined by the number of assets $n$, the number of considered time steps $m$, and the number of chosen asset units $k$. The risk-affinity/aversion parameter $\lambda$, the unit limit on asset selection $B$, and the capital bound $C$ have a strong influence on the solvability of the instances.
Moreover, we generated randomly perturbed instances based on real-world S&P 500 data from January 1, 2024, to May 31, 2024. The generated data is sampled independently at each time-step using the $t$-distribution with $\nu=5$ scaled to the mean and covariance of the stocks given by the original data with a rolling window of 30 days. Outliers are randomly added to the dataset. We always consider the $n$ biggest companies according to their market cap as of January 1, 2025. Table 5(right) gives an overview of the instances in the problem class, with the instances on top being randomly generated and the instances on the bottom representing real-world data.
@@ 4.6.4 Classical Baseline
Figure 11: (Portfolio) Runtimes (solid lines) and MIP Gap (dashed lines) for Gurobi with a time limit of 7200s for different values of $\lambda$ averaged over three randomly generated similar instances.
@@
0.000001 -1546160.82 -1535066.00 0.01 3644.56 -1540456 221.95 0.00
0.00001 -1522435.51 -1412403.50 0.08 3603.42 -1417302 245.25 0.00
0.00005 -1204575.81 -1185652.00 0.02 3601.04 -1186606 285.51 0.00
0.0001 -1104183.76 -1077731.50 0.02 3629.39 -1074430 296.73 0.00
0.0005 -1072027.81 -887057.50 0.21 3603.17 -826214 300.55 0.07
0.001 -1223385.04 -832785.50 0.47 3600.92 -711553 300.07 0.15
0.01 -5982818.73 136536.00 44.82 3600.75 -36277 295.94 1.27
Table 6: (Portfolio) Comparison of (classical) BQP vs.(quantum-ready) QUBO formulation of a portfolio optimization problem–a200_t10_s00 instance.
The tests are conducted by running the BQP model with Gurobi 11.0.0 on an AMD EPYC 7542 processor using 8 threads with a timeout of 3600s and using the QUBO model with ABS2 on a system with an A100 SXM4 80 GB GPU with a timeout of 300s. An overview for different parameters up to $n=400$ assets can be seen in Figure 11. With increasing $\lambda$, the running time and the upper bound on the optimality gap (MIP Gap) also increase. Table 6 shows execution details on the instance for $n=200$, $m=10$, $k=3$, $B=50$, $C=10$: All the objective function values shown are negative because we are minimizing the risk, and therefore, the profit is subtracted in the objective function. As may be expected, when the risk adversity increases, the profit decreases, as well as the gap between the lower and upper bound of the solution. The computational performance varies depending on the risk aversion parameter $\lambda$. For $\lambda=0$, the risk term is ignored, and only the profit is optimized. This can be quickly solved to optimality. In the risk-averse case with a high $\lambda=0.01$, the opposite happens. The lowest risk strategy is to take the risk-free interest rate $\nu=0.01$ and receive a profit of $C\cdot m\cdot\nu=1000$. The interesting cases are $\lambda=10^{-4}\ldots 10^{-6}$. Here, we can observe that ABS2 with the QUBO formulation finds equal or even better solutions faster.
Figure 12 shows an in-depth analysis for a problem instance with $n=200$, $m=10$, $k=3$, $B=50$, $C=10$. More specifically, the Figure shows how the objective value changes over time for $\lambda=0.001$. The same system was used for Gurobi and two A40 RTX 48 GB for ABS2. Although BQP achieves a better final objective value, the QUBO approach reaches good solutions in a substantially shorter time–without developing specialized heuristics.
Figure 12: (Portfolio) The plot illustrates the best found objective values at every step of the solving for different risk aversion parameters with Abs2 and Gurobi solvers for one instance with $\lambda=0.001$ and 200 assets. While ABS2 finds good solutions faster, Gurobi eventually achieves better results.
@@ 4.7.1 Background
Why is it interesting? The unweighted maximum independent set problem is a classic NP-hard mathematical problem. Starting at problem sizes of several hundred variables, there are known instances that are hard to solve to proven optimality with existing methods [211], and which are difficult to tackle even heuristically. An interesting property of this problem class is that it is well suited for translation into relatively sparse QUBOs–often with small coefficients.
@@ 4.7.3 Instances
Any collection of graphs might be a suitable benchmark set for MIS. We aimed to collect some instances known to be challenging and added several smaller ones to enable the tracking of progress. We collected instances from various sources, in particular: Network Repository [196], Steiner Lib [132], OESIS Implementation Challenge [211], and self-generated graphs. Furthermore, we supply a program (see QOBLIB [193]) that takes a graph and an independent set and verifies that it is indeed a stable set in the graph.
@@ 4.7.4 Classical Baseline
We provide formulations as binary linear programs and as QUBO in the Independent Set model section in QOBLIB [193]. These can be solved by any suitable solver. We provide baseline results for the binary linear program using Gurobi 11.0.0 run on an AMD EPYC 7542 32-core processor using 64 threads with a timeout of 7200 seconds. The baseline results for the QUBO formulation are computed using ABS2 as a heuristic run on four Tesla V100-SXM2-16GB with a timeout of 600 seconds. Moreover, we tested KaMIS (ReduMIS) [153, 111] on a system with an AMD EPYC 7542 using 32 threads. Corresponding runtime, TTS, and absolute gaps to the optimal/best-known solutions can be found in Figure 13. Exact data and instance names can be obtained from Table 7.
@@
Figure 13: (Independent Set) Runtimes, time-to-solution, and absolute gap to optimal/best-known solution for the problem sorted by Gurobi’s runtime. Only gaps $>0$ are plotted.
Gurobi KaMIS Abs2
# Name Nodes Edges Best RT (s) TTS (s) Gap TTS (s) Gap TTS (s) Gap Source
1 aves-sparrow-social 52 516 13 0.0 0.0 0 0.1 0 0.0 0 [196]
2 farm 17 39 10 0.0 0.0 0 0.0 0 0.0 0 [196]
3 football 35 118 16 0.0 0.0 0 0.0 0 0.0 0 [196]
4 chesapeake 39 170 17 0.0 0.0 0 0.1 0 0.0 0 [196]
5 karate 34 78 20 0.0 0.0 0 0.0 0 0.0 0 [196]
6 mammalia-kangaroo-interactions 17 91 4 0.0 0.0 0 0.0 0 0.0 0 [196]
7 ibm32 32 94 13 0.0 0.0 0 0.0 0 0.0 0 [196]
8 es60fst01 123 159 60 0.0 0.0 0 0.0 0 0.0 0 [132]
9 insecta-ant-colony1-day38 56 1134 6 0.0 0.0 0 0.2 0 0.0 0 [196]
10 es60fst03 113 142 55 0.0 0.0 0 0.0 0 0.0 0 [132]
11 es60fst02 186 280 88 0.0 0.0 0 0.0 0 0.0 0 [132]
12 es60fst04 162 238 78 0.0 0.0 0 0.0 0 0.0 0 [132]
13 sloane_1dc_64 64 543 10 0.0 0.0 0 0.1 0 0.0 0 [211]
14 sloane_1dc_128 128 1471 16 0.0 0.0 0 6.1 0 0.0 0 [211]
15 sloane_2dc_128 128 5173 5 0.1 0.0 0 0.6 0 0.0 0 [211]
16 sloane_1zc_128 128 2240 18 0.1 0.0 0 3.7 0 0.0 0 [211]
17 gen200_p0-9_44 200 1990 44 0.2 0.0 0 4.7 0 0.0 0 [196]
18 insecta-ant-colony3-day09 160 8883 9 0.3 0.0 0 0.9 0 0.0 0 [196]
19 C125 125 787 34 0.8 0.0 0 0.2 0 0.0 0 [196]
20 keller4 171 5100 11 1.3 0.0 0 0.7 0 0.0 0 [196]
21 brock200_2 200 10024 12 8.0 5.0 0 29.8 0 0.1 0 [196]
22 socfb-haverford76 1446 59589 282 978.7 36.0 0 20.1 0 3.6 0 [196]
23 keller6 3361 1026582 59 7200.0 3767.0 9 264.9 0 147.4 0 [196]
24 brock400_1 400 20077 27 7200.0 578.0 2 30.3 2 14.0 0 [196]
25 R_1000_005_1 1000 24670 117 7200.0 199.0 11 21.0 1 4.9 2 [198]
26 R_500_005_1 500 6256 91 7200.1 19.0 7 14.1 0 0.1 1 [198]
27 C500 500 12418 57 7200.1 567.0 2 12.0 0 0.3 0 [196]
28 socfb-trinity100 2613 111996 499 7200.1 297.0 1 21.3 0 42.6 1 [196]
29 hamming10-4 1024 89600 40 7200.1 235.0 0 64.1 0 0.1 0 [196]
30 frb100-40 4000 572774 100 7200.1 3516.0 11 327.6 4 47.3 9 [196]
31 sorrell7 2048 39424 198 7200.1 2529.0 13 36.2 0 191.3 10 [84]
32 p_hat1500-3 1500 277006 94 7200.1 1315.0 0 139.6 0 0.1 0 [196]
33 C4000-5 4000 3997732 18 7200.1 7200.1 6 4372.9 1 1.6 0 [196]
34 brock800_1 800 112095 23 7200.1 239.0 3 123.7 2 0.2 2 [196]
35 frb59-26-2 1534 126163 59 7200.2 241.0 4 74.0 1 64.4 3 [196]
36 sorrell4 2048 504451 24 7200.2 267.0 1 356.8 0 9.8 0 [196]
37 frb53-24-1 1272 94227 53 7200.5 630.0 3 41.8 2 12.6 2 [196]
38 frb50-23-3 1150 81068 50 7201.1 934.0 2 64.3 1 3.3 2 [196]
39 frb45-21-3 945 58245 45 7202.2 114.0 2 33.3 1 0.6 2 [196]
40 p_hat1500-1 1500 839327 12 7202.9 134.0 1 1447.8 0 0.6 0 [196]
@@
Table 7: (Independent Set) Best solution, Runtimes, TTS, and absolute gaps to best found (some even optimal) solutions for Gurobi, KaMIS (ReduMIS), and Abs2. Notably, only Gurobi checks optimality. Abs2 and KaMIS run as a heuristic. Please note that we do not report any runtimes for KaMIS or ABS2, since the stopping criteria for these heuristic solvers vary. The bold digits indicate optimal solutions in the best column.
@@ 4.8.1 Background
Network design models have wide applications in telecommunications, transportation planning [80, 178], and power networks [206]. One common formulation for this problem is to use an MIP formulation [21]. Given an $n\times n$ matrix $T$, where each entry $t_{ij}$ represents the traffic to be routed between vertex $i$ and $j$, and an integer $p>0$, construct a simple directed graph $D=(V,A)$ with vertex set $V=\{1,\ldots,n\}$ and a subset of selected edges $S^{*}\subseteq A$, where each vertex has indegree and outdegree equal to $p$. Furthermore, all the demands are routed such that the maximum load on any single edge of the network is minimized. The load on an edge is the sum of all the traffic that passes through that edge.
Why is it interesting? The min-max objective function underlying network design models is linear, and the problem often has a sparse structure. While it is simple to find feasible solutions, finding optimal solutions is much more difficult. The particular problem presented in this section has been withstanding all attempts to solve it to proven optimality for the last 30 years in its 24-node size. Furthermore, network design models have wide applications in telecommunications and transportation planning [80, 178].
Although we will focus on the mathematical formulation as a multicommodity-directed model, many other formulations have been used in network design, depending on the problem context. In [82, 6], the authors provide surveys of alternative formulations. In the standard formulation, both flow decisions variables and design decisions variables are modeled explicitly [21, 17, 82]. In this case, the linear programming relaxations of multicommodity flow formulations provide lower bounds that are normally weak [82, 6]. To overcome this issue, in [6], the authors remove the flow variables as explicit decisions and embed them within the design variables. Furthermore, in [82, 6] authors also provide a survey of different methods that have been used for network design. For example, in [174], the authors apply Benders decomposition and demonstrate the effectiveness of variable elimination preprocessing and a dual ascent procedure to accelerate the decomposition algorithm. In [21], authors report on computational experience with a cutting plane algorithm for the above formulation. In [32], they use the column generation technique to obtain near-optimal solutions.

### v2

@@ 4.6.3 Instances
Table 5 (left) outlines the parameters that are used in our experiments. The parameter choices are based on established literature [93, 57, 226]. Using these parameters, we generated 4 instances each (1 containing the original data and 3 with randomly perturbed data) for 10, 50, 200, and 400 possible assets with two different time horizons: 10 days and 15 days. We used real S&P 500 data from January 1, 2024, to May 31, 2024, for the asset prices $p_{it}$ and the corresponding covariances $\sigma_{ijt}$. The size of the instances is defined by the number of assets $n$, the number of considered time steps $m$, and the number of chosen asset units $k$. The risk-affinity/aversion parameter $\lambda$, the unit limit on asset selection $B$, and the capital bound $C$ have a strong influence on the solvability of the instances. We would like to highlight that the repository includes instances with $\lambda=0$, which represents an edge-case setting that lacks operational meaning but serves as a simplified mathematical representation.
Moreover, we generated randomly perturbed instances based on real-world S&P 500 data from January 1, 2024, to May 31, 2024. The generated data is sampled independently at each time step using the $t$-distribution with $\nu=5$, scaled to the mean and covariance of the stocks as given by the original data, with a rolling window of 30 days. Outliers are randomly added to the dataset. We always consider the $n$ largest companies, ranked by their market capitalization as of January 1, 2025. Table 5(right) gives an overview of the instances in the problem class, with the instances on top being randomly generated and the instances on the bottom representing real-world data.
@@ 4.6.4 Classical Baseline
Figure 11: (Portfolio) Runtimes (solid lines) and MIP Gap (dashed lines) for Gurobi with a time limit of 3600s for different values of $\lambda$ averaged over three randomly generated similar instances.
@@
0 -879572 -879572 0.00 0.4 -879572 90.0 0.00
0.000001 -872141 -872055 0.01 398.7 -872055 79.1 0.00
0.00001 -887265.36 -819899 8.21 3600.1 -819899 30.2 0.00
0.00005 -692341.57 -680124 1.80 3600.2 -680124 50.7 0.00
0.0001 -604495.30 -586823 3.01 3600.2 -586823 206.7 0.00
0.0005 -424863.29 -386990 9.80 3600.1 -384862 1497.6 0.55
0.001 -367893.74 -314736 16.90 3600.2 -299298 741.9 4.91
0.01 -141194.79 -437920 3124.21 3661.8 -2825 1510.8 99.35
Table 6: (Portfolio) Comparison of (classical) BQP using Gurobi 12.0.1 with a tolerance of 1e-4 vs.(quantum-ready) QUBO formulation of a portfolio optimization problem–a050_t15_s00 instance. The ABS2 Gap is the relative gap between the ABS2 incumbent and the gurobi incumbent.
The tests are conducted by running the BQP model with Gurobi 12.0.1 on an Intel Xeon Gold 6338 CPU @ 2.00GHz processor using 128 threads with a timeout of 3600s and using the QUBO model with ABS2 on a system with four A40 SXM4 48 GB GPUs with a timeout of 3600s. An overview of different parameters up to $n=50$ assets can be seen in Figure 11. With increasing $\lambda$, the running time and the upper bound on the optimality gap (MIP Gap) also increase. Table 6 shows execution details on the instance for $n=50$, $m=15$, $k=3$, $B=20$, $C=10$: All the objective function values shown are negative because we are minimizing the risk, and therefore, the profit is subtracted in the objective function. As may be expected, when the risk adversity increases, the profit decreases, as well as the gap between the lower and upper bounds of the solution. The computational performance varies depending on the risk aversion parameter $\lambda$. For $\lambda=0$, the risk term is ignored, and only the profit is optimized. This can be quickly solved to optimality. In the risk-averse case with a high $\lambda=0.01$, the opposite happens. The lowest risk strategy is to take the risk-free interest rate $\nu=0.01$ and receive a profit of $C\cdot m\cdot\nu=1000$.
Figure 12 shows an in-depth analysis for a problem instance with $n=50$, $m=15$, $k=3$, $B=20$, $C=10$. More specifically, the Figure shows how the objective value changes over time for $\lambda=0.0005$. The same system was used for Gurobi and ABS2. Although BQP achieves a better final objective value, the QUBO approach reaches good solutions in a substantially shorter time–without developing specialized heuristics.
Figure 12: (Portfolio) The plot illustrates the best found objective values at every step of the solving for different risk aversion parameters with Abs2 and Gurobi solvers for one instance with $\lambda=0.0005$ and 50 assets. While ABS2 finds good solutions faster, Gurobi eventually achieves better results.
@@ 4.7.1 Background
Why is it interesting? The unweighted maximum independent set problem is a classic NP-hard mathematical problem. Starting at problem sizes of several hundred variables, there are known instances that are hard to solve to proven optimality with existing methods [220], and which are difficult to tackle even heuristically. An interesting property of this problem class is that it is well-suited for translation into relatively sparse QUBOs–often with small coefficients.
@@ 4.7.3 Instances
Any collection of graphs might be a suitable benchmark set for MIS. We aimed to collect some instances known to be challenging and added several smaller ones to facilitate progress tracking. We collected instances from various sources, in particular: Network Repository [205], SteinLib [139], OESIS Implementation Challenge [220], and self-generated graphs. Furthermore, we provide a program that takes a graph and an independent set and verifies that it is indeed a stable set in the graph.
@@ 4.7.4 Classical Baseline
We provide formulations as binary linear programs and as QUBO. These can be solved by any suitable solver. We provide baseline results for the binary linear program using Gurobi 12.0.1 run on an Intel Xeon Gold 6338 CPU @ 2.00 GHz processor using 128 threads with a timeout of 7200 seconds. The baseline results for the QUBO formulation are computed using ABS2 as a heuristic run on four Nvidia A40 48GB with a timeout of 7200 seconds. Corresponding runtime, TTS, and absolute gaps to the optimal/best-known solutions can be found in Figure 13. Exact data and instance names can be obtained from Table 7.
@@
Figure 13: (Independent Set) Runtimes, time-to-solution, and absolute gap to optimal/best-known solution for the problem sorted by Gurobi’s runtime. Only gaps >0>0 are plotted.
Gurobi Abs2
# Name Nodes Edges Best RT (s) TTS (s) Gap TTS (s) Gap Source
4 farm 17 39 10 0.0 0.0 0 0.0 0 [205]
2 mammalia-kangaroo-interactions 17 91 4 0.0 0.0 0 0.0 0 [205]
3 johnson8-2-4 28 210 7 0.0 0.0 0 0.0 0 [205]
8 ibm32 32 90 13 0.0 0.0 0 0.0 0 [205]
5 karate 34 78 20 0.0 0.0 0 0.0 0 [205]
1 football 35 118 16 0.0 0.0 0 0.0 0 [205]
7 chesapeake 39 170 17 0.0 0.0 0 0.0 0 [205]
16 MANN-a9 45 918 3 0.1 0.0 0 0.0 0 [205]
10 aves-sparrow-social 52 454 13 0.0 0.0 0 0.0 0 [205]
11 insecta-ant-colony1-day38 56 1134 6 0.0 0.0 0 0.0 0 [205]
6 sloane_1dc_64 64 543 10 0.0 0.0 0 0.0 0 [205]
23 hamming6-4 64 704 12 0.2 0.0 0 0.0 0 [220]
12 hamming6-2 64 1824 2 0.0 0.0 0 0.0 0 [205]
20 johnson8-4-4 70 1855 5 0.1 0.0 0 0.0 0 [139]
13 es60fst03 113 142 55 0.0 0.0 0 0.0 0 [205]
19 johnson16-2-4 120 5460 15 0.1 0.0 0 0.0 0 [205]
15 es60fst01 123 159 60 0.0 0.0 0 0.0 0 [205]
26 C125-9 125 787 34 1.0 0.0 0 0.0 0 [220]
24 sloane_1zc_128 128 1120 18 0.2 0.0 0 0.0 0 [220]
22 sloane_1dc_128 128 1471 16 0.1 0.0 0 0.0 0 [139]
18 sloane_2dc_128 128 5173 5 0.1 0.0 0 0.0 0 [139]
25 insecta-ant-colony3-day09 160 8883 9 0.2 0.0 0 0.0 0 [220]
9 es60fst04 162 238 78 0.0 0.0 0 0.1 0 [205]
27 keller4 171 5100 11 1.7 0.0 0 0.0 0 [205]
14 es60fst02 186 280 88 0.0 0.0 0 0.0 0 [205]
17 c-fat200-1 200 1534 18 0.1 0.0 0 0.0 0 [205]
21 gen200_p0-9_44 200 1990 44 0.1 0.0 0 0.0 0 [139]
29 brock200-2 200 10024 12 6.0 3.0 0 0.0 0 [205]
31 brock200-3 200 12048 9 7.6 0.0 0 0.0 0 [205]
30 brock200-4 200 13089 8 6.3 0.0 0 0.0 0 [205]
28 brock200-1 200 14834 6 5.6 0.0 0 0.0 0 [205]
37 brock400-1 400 20077 27 7200.1 1269.0 0 8.4 0 [205]
33 R_500_005_1 500 6256 91 7200.0 86.0 2 0.4 1 [205]
35 C500-9 500 12418 57 7200.1 20.0 3 7.5 0 [207]
45 brock800-1 800 112095 23 7200.3 395.0 3 1139.3 0 [205]
49 frb45-21-3 945 58245 45 7201.5 806.0 1 8.7 2 [205]
39 R_1000_005_1 1000 24670 117 7200.1 2859.0 9 169.6 2 [205]
41 hamming10-4 1024 89600 40 7200.2 108.0 0 0.1 0 [89]
42 frb50-23-3 1150 81068 50 7200.2 3226.0 2 252.3 1 [205]
40 frb53-24-1 1272 94227 53 7200.1 903.0 3 1023.5 2 [205]
32 socfb-haverford76 1446 59589 282 838.7 26.0 0 0.6 0 [205]
47 p_hat1500-3 1500 277006 94 7200.8 3915.0 0 0.2 0 [205]
48 p_hat1500-1 1500 839327 12 7201.4 523.0 1 2.4 0 [205]
46 frb59-26-2 1534 126163 59 7200.8 3980.0 3 1.9 3 [205]
34 sorrell7 2048 39424 198 7200.1 513.0 13 663.5 9 [205]
44 sorrell4 2048 504451 24 7200.2 257.0 1 1708.8 0 [205]
43 socfb-trinity100 2613 111996 499 7200.2 1891.0 1 1361.5 0 [205]
36 keller6 3361 1026582 59 7200.1 2377.0 18 290.9 0 [207]
50 frb100-40 4000 572774 100 7205.1 5695.0 23 4634.8 7 [205]
38 C4000-5 4000 3997732 18 7200.1 7200.1 6 167.0 0 [205]
@@
Table 7: (Independent Set) Best solution, Runtimes, TTS, and absolute gaps to best found (some even optimal) solutions for Gurobi 12.0.1 and Abs2. Notably, only Gurobi checks optimality. Abs2 runs as a heuristic. Please note that we do not report any runtimes for ABS2, since the stopping criteria for these heuristic solvers vary.
@@ 4.8.1 Background
Network design models have wide applications in telecommunications, transportation planning [85, 187], and power networks [215]. One common formulation for this problem is to use an MIP formulation [22]. Given an $n\times n$ matrix $T$, where each entry $t_{ij}$ represents the traffic to be routed between vertex $i$ and $j$, and an integer $p>0$, construct a simple directed graph $D=(V,A)$ with vertex set $V=\{1,\ldots,n\}$ and a subset of selected edges $S^{*}\subseteq A$, where each vertex has indegree and outdegree equal to $p$. Furthermore, all the demands are routed such that the maximum load on any single edge of the network is minimized. The load on an edge is the sum of all the traffic that passes through that edge. The problem is NP-hard [22].
Why is it interesting? The min-max objective function underlying network design models is linear, and the problem often has a sparse structure. While it is trivial to construct a feasible solution, finding optimal solutions is very difficult. The particular problem presented in this section has been withstanding all attempts to solve it to proven optimality for the last 30 years in its 24-node size. Furthermore, network design models have wide applications in telecommunications and transportation planning [85, 187].
Although we will focus on the mathematical formulation as a multicommodity-directed model, many other formulations have been used in network design, depending on the problem context. In [87, 6], the authors provide surveys of alternative formulations. In the standard formulation, both flow decision variables and design decision variables are modeled explicitly [22, 17, 87]. In this case, the linear programming relaxations of multicommodity flow formulations provide lower bounds that are normally weak [87, 6]. To overcome this issue, in [6], the authors remove the flow variables as explicit decisions and embed them within the design variables. Furthermore, in [87, 6] the authors also provide a survey of different methods that have been used for network design. For example, in [183], the authors apply Benders decomposition and demonstrate the effectiveness of variable elimination preprocessing and a dual ascent procedure to accelerate the decomposition algorithm. In [22], the authors report on computational experience with a cutting plane algorithm for the above formulation. In [36], they use the column generation technique to obtain near-optimal solutions.